Required information [The following information applies to the questions displayed below.] Summary information from the finan2A Prof Mar Ratio 2A Price Earn 2A Ret On 2A Tot Asset 2A Ret on Tot 2A Div Yield Req 2B Com Stock Turn Assets Ratio For both2A Ret On 2A Prof Mar 2A Ret on Tot 2A Tot Asset 2A Price Earn Req 2B 2A Div Yield Ratio Turn Assets Com Stock Ratio For both2A Price Earn 2A Ret on Tot 2A Prof Mar 2A Tot Asset 2A Ret On 2A Div Yield Req 2B Turn Com Stock Ratio Assets Ratio For both2A Prof Mar 2A Price Earn 2A Tot Asset 2A Ret on Tot 2A Ret On 2A Div Yield Req 2B Ratio Com Stock Ratio Turn Assets 2A Tot A2A Ret On 2A Price Earn Ratio 2A Prof Mar 2A Tot Asset 2A Ret on Tot 2A Div Yield Req 2B Com Stock Ratio Turn Assets Assuming2A Price Earn Ratio 2A Prof Mar 2A Ret On 2A Ret on Tot 2A Tot Asset 2A Div Yield Req 2B Ratio Turn Assets Com Stock Assuming2A Price Earn 2A Prof Mar 2A Tot Asset 2A Ret on Tot 2A Ret On 2A Div Yield Req 2B Com Stock Ratio Turn Assets Ratio Identify
Answer #1

2A profit margin ratio

Profit Margin Ratio

Company

Choose Numerator:

/

Choose Denominator:

=

Profit margin ratio

Net income

/

Net sales

=

Profit margin ratio

Barco

$180523

/

$800,000

=

22.6

%

Kyan

$198345

/

$882200

=

22.5

%

2A Total asset turnover

Total Asset Turnover

Company

Choose Numerator:

/

Choose Denominator:

=

Total Asset Turnover

Net sales

/

Average total assets

=

Total asset turnover

Barco

$800,000

/

$438470

=

1.8

times

Kyan

$882200

/

$468275

=

1.9

times

Average total assets = (beginning total assets+ending total assets)/2

Barco = (408000+468940)/2 = 438470

Kyan = (                392500+544050)/2 = 468275

2A Return on Total Assets

Return on Total Assets

Company

Choose Numerator:

/

Choose Denominator:

=

Return on Total Assets

Net income

/

Average total assets

=

Return on total assets

Barco

$180523

/

$438470

=

41.2

%

Kyan

$198345

/

$468275

=

42.4

%

2A Return on Common Stockholders’ Equity

Return On Common Stockholders' Equity

Company

Choose Numerator:

/

Choose Denominator

=

Return On Common Stockholders' Equity

Net income

-

Preferred dividends

/

Average common stockholders' equity

=

Return on common stockholders' equity

Barco

$180523

-

/

$300659

=

60.0

%

Kyan

$198345

-

/

$344742

=

57.5

%

Average common stockholders' equity = (beginning common stockholders' equity + ending common stockholders' equity)/2

Barco = (210000+78517+210000+102800)/2 = 300659

Kyan = (236000+103733+236000+113750)/2 = 344742

2A Price-Earnings Ratio

Price-Earnings Ratio

Company

Choose Numerator:

/

Choose Denominator:

=

Price-Earnings Ratio

Market price per common share

/

Earnings per share

=

Price-Earnings Ratio

Barco

$75

/

$4.30

=

17.4

times

Kyan

$75

/

$4.20

=

17.6

times

2A Dividend Yield

(f)

Dividend Yield

Company

Choose Numerator:

/

Choose Denominator:

=

Dividend Yield

Annual cash dividends per share

/

Market price per share

=

Dividend Yield

Barco

$3.72

/

$75

=

5.0

%

Kyan

$3.99

/

$75

=

5.3

%

Part 2 B

Investment in Kyan company is the better investment as its profitability ratios are high which states investment in it will be more profitable.